OpenStudy (anonymous):

if the limit as x approaches -3 of f(x)/x^2, find the limit as x approaches -3 of f(x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry, its supposed to say if the limit as x approaches -3 of f(x)/(x^2)=1, find...

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

i really don't know. i do not understand limits

OpenStudy (aum):

You can directly substitute x = -3 into the expression and figure out f(-3).

OpenStudy (anonymous):

what does the bottom (denominator) go to?

OpenStudy (anonymous):

so would i graph x/x^2 and see where it equals -3?

OpenStudy (anonymous):

the denominator goes to x(x)

OpenStudy (anonymous):

no... substitute in x = -3 into the bottom function. what do you get?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

\[\lim_{x \rightarrow -3}\frac{ f \left( x \right) }{ x^2}=\frac{ \lim_{x \rightarrow -3} f \left( x \right)}{ \lim_{x \rightarrow -3} x^2}=1 \Rightarrow \lim_{x \rightarrow -3} f \left( x \right)=\lim_{x \rightarrow -3} x^2\]

OpenStudy (anonymous):

learn your properties of limits!